# ChunkRAG: Novel LLM-Chunk Filtering Method for RAG Systems

### v2

@@ 3 Methodology
Figure 2: Methodology
Step 1: Semantic Chunking
@@
    Input Preparation: Start with a document $D$ and break it down into sentences using a sentence tokenizer (NLTK’s $sent\_tokenize$ function).
@@
Step 2: Vector Store Creation
@@
Step 3: Retriever Initialization
@@
Step 4: Query Rewriting
@@
Step 5: Initial Filtering
@@
Step 6: Advanced Relevance Scoring
@@
Step 7: Thresholding for Relevance
@@
    Set Dynamic Threshold: LLM analyzes the distribution of final scores and suggests an optimal threshold for chunk relevance.
@@
Step 8: Hybrid Retrieval and Re-Ranking
@@
    Dual Retrieval Methods: Combine BM25 and LLM-based retrieval methods to ensure a balanced retrieval strategy. BM25 retriever is implemented to capture keyword-based retrieval. BM25 retriever is combined using an ensemble approach with equal weights (0.5 each).
@@
    Re-Rank Chunks: Cohere’s reranking model (‘rerank-englishv3.0‘) is applied to rank the chunks with relevance to deal with Lost in the middle problem
Step 9: Answer Generation
@@
    Accuracy=Number of Correct AnswersTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Answers}}{\text{Total Number of Questions}}$
@@ 4 Experiments
We conducted experiments to extensively demonstrate ChunkRAG’s adaptability and its potential for generalizability across various generation tasks. However, due to computational resource constraints, our evaluation was primarily focused on the PopQA dataset.
@@ 4.1 Tasks and Datasets
ChunkRAG was evaluated on the PopQA dataset, which serves as the cornerstone of our experimental analysis. PopQA Mallen et al. (2023) is a benchmark dataset designed for short-form question answering. It comprises a diverse set of questions that require concise and accurate responses, making it an ideal testbed for assessing the performance of retrieval-augmented generation models like ChunkRAG.
To measure the effectiveness of ChunkRAG, we adopted accuracy as the evaluation metric, consistent with prior studies. This metric aligns with the conventions used in the evaluation of PopQA, ensuring that our results are comparable to existing research.
While our current experiments are limited to PopQA, ChunkRAG is architected with scalability in mind. Future evaluations may extend to additional datasets such as Biography Min et al. (2023) for long-form generation, PubHealth Zhang et al. (2023) for true/false question answering, and Arc-Challenge Bhakthavatsalam et al. (2021) for multiple-choice questions. These extensions will further validate ChunkRAG’s versatility across different types of generation tasks, contingent upon the availability of computational resources.
@@ 4.2 Baselines
4.2.1 Baselines Without Retrieval
We first evaluated several large language models (LLMs) that do not incorporate retrieval mechanisms. Among the public LLMs, we included LLaMA2-7B and LLaMA2-13B Touvron et al. (2023), known for their versatility across diverse natural language processing (NLP) tasks, and Alpaca-7B and Alpaca-13B Dubois et al. (2023), which are instruction-tuned models optimized for effectively following user prompts. Additionally, we assessed CoVE65B Dhuliawala et al. (2024), which introduces iterative engineering techniques aimed at enhancing the factual accuracy of generated content. For proprietary models, we included LLaMA2-chat13B, a conversational variant of LLaMA2 tailored for dialogue-based applications, and ChatGPT, OpenAI’s proprietary conversational agent renowned for its robust language understanding and generation capabilities.
4.2.2 Baselines With Retrieval
Standard Retrieval-Augmented Generation (RAG): To establish a baseline for retrieval-augmented methods, we evaluated standard RAG approaches. Specifically, we employed Standard RAG Lewis et al. (2020), which utilizes a retriever to fetch relevant documents based on the input query, subsequently feeding these documents into the language model to generate responses. For consistency, we utilized the same retriever mechanism as ChunkRAG to ensure a fair comparison. In addition to Standard RAG, we evaluated instruction-tuned LLMs with standard RAG, including LLaMA2-7B, LLaMA2-13B, and Alpaca-7B, Alpaca-13B, to assess the impact of instruction tuning in conjunction with retrieval augmentation.
Advanced Retrieval-Augmented Generation: To benchmark ChunkRAG against more sophisticated RAG-based methods, we included advanced models that incorporate additional strategies to enhance performance. SAIL Luo et al. (2023) enhances standard RAG by instruction-tuning the language model on Alpaca instruction-tuning data, inserting top retrieved documents before the instructions to provide contextual information. Self-RAG Asai et al. (2024) further refines RAG by incorporating reflection tokens labeled by GPT-4 within the instruction-tuning data, enabling the model to better utilize retrieved information. Additionally, we considered CRAG Your et al. (2024), a recent approach that augments standard RAG with corrective strategies to improve retrieval quality by addressing low-quality retrieval results. CRAG serves as a direct comparison to our proposed ChunkRAG, highlighting the effectiveness of our chunk filtering mechanism in enhancing retrieval-augmented generation. Furthermore, we evaluated retrieval-augmented baselines with private data, including Ret-ChatGPT and RetLLaMA-chat, which integrate retrieval mechanisms with ChatGPT and the conversational variant of LLaMA2, respectively.

### v3

@@ 3 Methodology
3.1 Step 1: Semantic Chunking
@@
    Input Preparation: Start with a document $D$ and break it down into sentences using a sentence tokenizer (e.g., NLTK’s sent_tokenize function).
@@
3.2 Step 2: Vector Store Creation
@@
3.3 Step 3: Retriever Initialization
@@
3.4 Step 4: Query Rewriting
@@
3.5 Step 5: Initial Filtering
@@
3.6 Step 6: Advanced Relevance Scoring
@@
3.7 Step 7: Thresholding for Relevance
@@
    Set Dynamic Threshold: The LLM analyzes the distribution of final scores and suggests an optimal threshold for chunk relevance.
@@
3.8 Step 8: Hybrid Retrieval and Re-Ranking
@@
    BM25 Retriever: In parallel, a BM25 retriever is implemented to capture keyword-based retrieval. The BM25 retriever is combined with the LLM-based retriever using an ensemble approach with equal weights (0.5 each), ensuring both semantic and keyword-based retrieval are balanced.
@@
    Re-ranking Chunks: Cohere’s reranking model (rerank-englishv3.0) is applied to rank the chunks by relevance to address the “Lost in the Middle” problem, ensuring the most relevant chunks are prioritized in the final retrieval set.
3.9 Step 9: Answer Generation
@@ 3.10 Step 10: Evaluation
    Accuracy=Number of Correct AnswersTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Answers}}{\text{Total Number of Questions}}$ (1)
@@ 4 Experiments
To evaluate the effectiveness of ChunkRAG in reducing irrelevance and hallucinations in retrieval-augmented generation, we conducted a series of experiments using the PopQA dataset. This section details the objectives, datasets, experimental design, baseline comparisons, and results.
4.1 Objectives
The primary objective of our experiments was to assess ChunkRAG’s performance in accurately retrieving relevant information and minimizing hallucinations, especially on tasks requiring high precision in short-form question answering. We hypothesized that the fine-grained, chunk-level filtering would improve retrieval accuracy compared to document-level approaches.
@@ 4.2 Tasks and Datasets
ChunkRAG was evaluated on the PopQA dataset Mallen et al. (2023), chosen for its alignment with our model’s goal of precise, fact-based retrieval. PopQA is a benchmark dataset designed for short-form question answering and contains diverse questions requiring concise, accurate responses. Its focus on factual precision makes it an ideal testbed for assessing the capabilities of retrieval-augmented generation (RAG) systems. While our primary experiments focused on PopQA, ChunkRAG is designed for scalability, and future evaluations could include additional datasets like Biography Min et al. (2023) for long-form generation, PubHealth Zhang et al. (2023) for true/false question answering, and Arc-Challenge Bhakthavatsalam et al. (2021) for multi-choice questions.
4.3 Experimental Setup
We conducted our experiments with computational constraints, implementing ChunkRAG on the PopQA dataset to evaluate its retrieval accuracy and response quality. The cosine similarity threshold for semantic chunking was set to $\theta=0.7$, marking chunk boundaries when similarity between consecutive sentences dropped below this level. Each chunk was limited to a maximum of 500 characters, ensuring focused and relevant segments.
In addition to the chunking process, we employed initial filtering via TF-IDF and cosine similarity for duplicate removal, followed by an advanced relevance scoring system incorporating a critic model. These settings were chosen based on preliminary analyses to optimize for relevance and accuracy.
@@ 4.4 Baselines
We compared ChunkRAG to several baselines, both with and without retrieval mechanisms:
Baselines Without Retrieval:
This group includes several large language models (LLMs) without retrieval mechanisms, such as LLaMA2-7B and Alpaca-7B Dubois et al. (2023), selected for their performance in various natural language processing tasks. Additionally, CoVE65B Dhuliawala et al. (2024) was included to measure factual accuracy improvements in non-retrieval LLMs.
Baselines With Retrieval:
For retrieval-augmented methods, we evaluated Standard RAG Lewis et al. (2020), Self-RAG Asai et al. (2024), and CRAG Your et al. (2024). CRAG serves as a direct comparison due to its corrective strategies in retrieval quality, though it lacks chunk-level filtering. The advanced models, such as Ret-ChatGPT and RetLLaMA-chat, were also included for proprietary comparisons.
4.5 Evaluation Metrics
The primary evaluation metric was accuracy, defined as the percentage of generated responses that exactly matched the ground-truth answers. We also assessed semantic accuracy, where responses semantically equivalent to the ground truth, as evaluated by an LLM-based similarity checker, were counted as correct. This approach ensured a comprehensive measure of both factual accuracy and contextual relevance.
Accuracy=Number of Correct AnswersTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Answers}}{\text{Total Number of Questions}}$ (2)
4.6 Results and Analysis
As shown in Table 1, ChunkRAG achieved an accuracy of 64.9 on PopQA, surpassing all baselines in the same category. Compared to CRAG, which achieved an accuracy of 54.9, our model shows a performance improvement of 10 percentage points. This increase is significant, especially for multi-step reasoning tasks, where errors can compound across steps.
The enhanced performance is primarily due to ChunkRAG’s chunk-level filtering, which reduces the inclusion of irrelevant or weakly related information. By focusing on semantically relevant chunks, the generation of factually accurate responses was substantially improved. Furthermore, our self-reflective scoring mechanism reduced retrieval errors by providing a finer relevance assessment at the chunk level.
Method PopQA Accuracy (%)
LLaMA2-13B 20.0
Ret-LLaMA2-13B 51.8
ChatGPT 29.3
Ret-ChatGPT 50.8
LLaMA2-7B 14.7
Alpaca-7B 23.6
Standard RAG 50.5
CRAG 54.9
Self-RAG 50.5
ChunkRAG (Ours) 64.9
Table 1: Performance Comparison Across Methods on PopQA Dataset
4.7 Observations and Insights
Our experiments demonstrated that chunk-level filtering led to a notable improvement in response accuracy and relevance. By dividing text into semantically coherent chunks, ChunkRAG was able to reduce irrelevant or tangential information, thus enhancing factual accuracy. The LLM’s self-reflective scoring system further contributed to error reduction by refining chunk relevance.
Future work will explore the scalability of ChunkRAG on additional datasets, including Biography, PubHealth, and Arc-Challenge, to validate its versatility. Addressing computational efficiency is also planned to enable broader applications of ChunkRAG.